An experiment consists of drawing a total of 8 cards from a deck (each time a single card is
drawn, then replaced into the deck, and the deck is shuffled before the next card is drawn).
(a) Find the probability of getting exactly 2 hearts.
(b) Find the probability of getting at most 2 hearts.
Please include steps in your solution? :)

Answer by stanbon(75887)

An experiment consists of drawing a total of 8 cards from a deck (each time a single card is drawn, then replaced into the deck, and the deck is shuffled before the next card is drawn).
(a) Find the probability of getting exactly 2 hearts.
Binomial Problem with n = 8 and p = 1/4
P(x = 2) = 8C2(1/4)^2(3/4)^6 = 0.3115
------------------------------------------------

(b) Find the probability of getting at most 2 hearts.
P(0<= x <=2) = binomcdf(8,1/4,2) = 0.6785
===================
Cheers,
Stan H.